# Model Synthesis for Zero-Shot Model Attribution

### v2

@@ III-A Preliminary Analysis of Model Fingerprint
Different Generative Blocks. The discussion above focuses on the influence of basic components. Moving forward, it is important to consider how different generative blocks throughout the entire generation process contribute to the final spectral pattern on the generate image. Generative models typically consist of multiple generative blocks, with upsampling layers playing a crucial role as connectors between these blocks, doubling the resolution of the input feature maps. As mentioned above, commonly used upsampling layers, such as bilinear and nearest-neighbor interpolation, inherently possess a low-pass filtering effect. This property leads them to attenuate high-frequency components from prior layers. Given that existing research [6] suggesting that model fingerprints predominantly reside in the high-frequency components, it becomes apparent that the later blocks in a model’s architecture could more significantly influence the model fingerprints on the output image as the high-frequency patterns from earlier layers could be filtered out. To empirically test this hypothesis, we designed an experiment depicted in Figure 7, consisting of three key steps:
@@
Figure 7(d,e) shows the confusion matrices under Testing I and Testing II of two sets of experiments: one set on two ProGAN models utilizing nearest-neighbor upsampling (d), and the other set on two ProGAN models employing bilinear-upsampling (e). As shown in the Figure, under Testing I, the classifier achieves high accuracy, indicated by the high diagonal values in the confusion matrix. However, after swapping the last few blocks in Testing II, the confusion matrix reveals a notable reversal in model attribution results. The shifted results indicate that most distinguishable spectrum patterns used to differentiate the models are predominantly from the models’ last few blocks. And the patterns left by earlier blocks might be less pronounced as they progress through the generative process.
In summary, different generative models exhibit distinct spectral patterns due to both architectural and parameter-related factors, including the types of upsampling and activation functions employed, normalization techniques, and the parameters of convolution layers. Notably, our observations indicate that within the architecture of a large generative model, the final few blocks tend to have a more profound influence on the distinguishable spectral patterns of the output images.
@@ III-B Model Synthesis Strategy
Overall, our model synthesis strategy is illustrated in Figure 8, the structure of the synthetic model could be viewed as a shallow auto-encoder, composed of $K$ downsampling blocks and $K$ upsampling blocks. Based on the above discussion, $K$ can selected as a small value, such as one or two. Each downsampling block use a fixed architecture, with a pooling layer to downscale the input resolution by half, and two convolutional layers to increase the feature dimension. The feature output from the downsampling blocks is then sent into the upsampling blocks, which share common components with the generative blocks in actual generative models and vary in many architecture options. The architecture of these blocks is defined by the options $\{U,A,N,L,S\}$.
@@
[…]
@@
We recognize that speed is key factor in model synthesis. In our approach, we simplify the objective to focus on reconstruction, where the generative neural network is specifically tasked with minimizing the reconstruction loss of input images. This method is not only straightforward to implement but also allows for rapid training. We leave the exploration of other training approaches to future work. We set a constraint on the minimum reconstruction residual to $\eta=0.005$ to limit the amount of artifacts in the output images. Ultimately, the total number of synthetic models generated is 5760, allowing for a wide simulation of a wide range of potential generative models with diverse architectures and parameters.
@@
We measured the time required to train these models. For models with $K=1$, the averaged required time per model is 53 seconds, while those with $K=2$ required 113 seconds each. In total, training all the synthetic models takes 133 GPU hours on a single 3090 GPU. This is considerably more efficient than the time investment needed for training state-of-the-art generative models. For example, it takes 14 days 22 hours on single V100 GPUs to train a StyleGAN [41] model that generate with a 256$\times$256 resolution [^4] https://github.com/NVlabs/stylegan.
III-C Training the Fingerprint Extractor
@@
$I_{k}=M_{k}(I_{\text{real}})$
These generated images, now embedded with model-specific fingerprints, are subsequently processed by a denoiser by [37]. The primary function of this denoiser is to strip away low-frequency semantic content, thereby revealing the underlying high-frequency fingerprint patterns inherent in the images. The extracted noises are then transformed using a Discrete Fourier Transform to facilitate the extraction of their spectrum properties. The spectrum obtained from this transformation are fed into the fingerprint extractor, which is built upon the ResNet50 architecture.
$h_{k}$ $=f_{\operatorname{fp\_extractor}}\left({\mathcal{F}}(I_{k}^{\prime})\right)$
@@
To enhance the model’s performance, we combine classification and metric loss to enhance the discrimination of learned fingerprint embedding. The classification head employs a multi-class cross-entropy (CE) loss function, which enable the fingerprint extractor to accurately classify the originating synthetic model of each input image. The projection head is followed by a triplet loss, which encourages generated images from the same generator to be as similar as possible in the feature space while promoting a larger dissimilarity between images from different generators.
$\mathcal{L}=\mathcal{L}_{\text{CE}}(h_{k},y_{k})+\mathcal{L}_{\text{Triplet}}(h_{a},h_{p},h_{n})$
where $y_{k}$ is the true label for $k$-th synthetic model. $h_{a}$, $h_{p}$, and $h_{n}$ are the anchor, positive, and negative fingerprint embedding from a triplet set.
@@ IV-A Experimental Details
Testing Models. Although our analysis and synthesis strategy primarily focuses on CNN-based generative models, our evaluation includes a broader scope of models as shown in Table I. In addition to CNN-based models, we also assess the generalization performance on Transformer-based models, which share many common components with CNN-based models. And for CNN-based category, we include a variety of models such as CNN, VAE, Flow, and Diffusion models. All models in the CNN-based and Transformer-based category are unconditional models, which are trained on the dataset in the “Image Source” column with different resolutions. Furthermore, the last row of the table encompasses state-of-the-art Text2Image models such as Stable Diffusion v1.5, Stable Diffusion v2.1, Stable Diffusion XL, Glide, DalleE-mini, DalleE-2, and DalleE-3. These models are not restricted to specific domains, and diverse in architecture type, training dataset, and resolution.
For the ablation study in Section IV-C, our fingerprint extractor is purely trained on synthetic models, all models in the table are unseen. For the comparisons with existing methods in Section IV-D, certain models highlighted in gray are used for training existing methods, and a fine-tuned version of our method pre-trained on synthetic models.
TABLE I: Models highlighted in gray are considered classical models and are used to train the existing methods being compared. The remaining models are categorized as emerging models and have not been seen by the existing methods.
@@
Experimental Setup. Similar to the face identification and verification problems encountered in the field of face recognition, we consider two model attribution problems, the 1:1 model verification and 1:N model identification problem, as shown in Figure 10. Settings and experimental setups for the two problems are illustrated below:
@@
  • Model Verification: the 1:1 model verification problem aims to verify whether two generated images are from the same model or not. Model verification is performed by comparing the similarity of the extracted fingerprint. To assess the performance, we generate 1,000 images for each model under test. Then we randomly selected 10,000 pairs out of these images, which consist of 5,000 negative pairs (images from different models) and 5,000 positive pairs (images from the same model). The evaluation employs metrics commonly used in face verification tasks: accuracy and the Area Under the Receiver Operating Characteristic Curve (AUC). Accuracy measures the correct verification of pairs as originating from the same or different models. We report the best accuracy across all similarity thresholds to assess the overall effectiveness of the verification system. The AUC provides a comprehensive measure of performance across all possible threshold levels, reflecting the trade-off between the true positive rate and the false positive rate.
@@
  • Model Identification: the 1:N model identification problem aims to identify the specific generative model used to create a given image among N models. To assess the performance of model identification, we generate 1,000 images for each model under test. Out of these, 10 images are sampled to create the fingerprint for each model, while the rest serve as test samples. Each model’s fingerprint is collected to form the model gallery. We compare the test sample with each fingerprint in the gallery and classify it to the class with the closest fingerprint distance. Performance is evaluated using classification accuracy and the F1 score.
@@
Compared methods. To compare with existing methods, we derive two variants of our method: 1) Ours (Only-Syn), which is only trained on synthetic models without having seen any real models in Table I. 2) Ours (Fine-tune), which fine-tune the fingerprint extractor trained on synthetic models for one epoch using samples of models highlighted in gray in the table.
@@
We compare against five existing model attribution methods. They are Marra et al. [5], Yu et al. [6], DNA-Det [8], RepMIX [9], and POSE [10]. For the training-based methods among these, we train them on models in gray in Table I. Then we extract the layers before the classification head as the fingerprint extractor. This adaptation is implemented to evaluate the model identification and verification problems.
IV-B Effectiveness of Model Synthesis (EQ1)
The generalization capability of our trained fingerprint extractor is built upon synthetic models. Before evaluating the fingerprint extractor, we first conduct evaluations on the synthetic models in two crucial dimensions: 1) Fidelity, which assesses the extent how well the synthetic models replicate the spectrum/fingerprint patterns of real models. 2) Diversity, which examine whether the synthetic models exhibit a wide range of spectrum patterns, rather than merely replicating a limited set.
Models All w/o K w/o L w/o S
FFD 3.30 3.57 3.41 3.51
Models w/o U w/o A w/o N w/o seed
FFD 9.81 4.46 3.02 3.39
Fidelity. Currently, there is no off-the-shelf tool to quantitatively measure the gap between synthetic and real models in fingerprint distributions. To address this gap, we introduce a new metric called Frechet Frequency Distance (FFD), inspired by the Frechet Inception Distance (FID) [62] commonly used to assess the quality of images from generative models. The FFD is calculated through two primary steps:
  • (Step I) Spectrum Pattern Representation: We extract the spectrum pattern of generated images to represent the fingerprint pattern of each synthesized and real models. As depicted in Figure 11(a), for each model, we initially generate $N$ (e.g., 100) images, and then, following [31], we enhance the spectrum patterns by applying a noise extractor [37] to remove semantic contents from these images. The resultant noise images are then used to compute a reduced 1D power spectrum, as specified in Equation 1. We calculate the averaged spectrum from these $N$ images to serve as the spectrum pattern representation for each model.
  • (Step II) Frechet Distance Calculation: As illustrated in Figure 11(b), we first compute the mean and covariance of the spectrum representations for real models in Table I, and synthetic models in Section 8. The Fréchet distance is then calculated between these distributions. Specifically, for the mean $\mu_{r}$ and covariance $\Sigma_{r}$ of real distribution, and the mean $\mu_{s}$ and covariance $\Sigma_{s}$ of synthetic distribution, their distance is given by:
    $\text{FFD}=\left\|\mu_{r}-\mu_{s}\right\|^{2}+\operatorname{Tr}\left(\Sigma_{r}+\Sigma_{s}-2\left(\Sigma_{r}\Sigma_{s}\right)^{1/2}\right),$
Figure 12: Compare the relative proportions of volume against the total volume of all synthesized models when reducing each synthesis option.
TABLE III: Ablation study on the architecture of synthetic models.
@@ IV-B Effectiveness of Model Synthesis (EQ1)
In Table II, we evaluate the Frechet Frequency Distance of synthetic models against real models, varying synthesis choices. The variations in the architecture include: ”w/o K” and ”w/o L,” indicating the construction with a single down/up block and a single layer, respectively; ”w/o S,” using only one type of sequence — normalization and activation after convolution; ”w/o U,” utilizing only bilinear interpolation for upsampling; and ”w/o A” and ”w/o N,” which denote the absence of activation and normalization layers, respectively. We also consider not using diverse seeds for each architecture, denoted as ”w/o seed”. The results displayed in the table indicate that reducing most synthesis options tends to increase the Frechet Frequency Distance (FFD), pointing to a poorer alignment between the distributions. Notably, eliminating the diversity in activation functions and upsampling methods results in the most significant increases in FFD. This may be because these factors, when varied, typically enhance the spectrum resemblance to real distributions.
Diversity. Although the FFD score effectively compares the fingerprint resemblance to the real models in Table I, the limited size of the real model dataset might prevent it from fully capturing the diversity of the synthetic data, potentially resulting in a bias toward the limited variety seen in the real dataset. Therefore, we propose using an additional metric, the Volume of Confidence Ellipsoid [63], to better assess the extent to which the synthetic models occupy the spectrum representation space. The volume is calculated as follows:
$\text{Volume}=\frac{(2\pi)^{n/2}}{\Gamma\left(\frac{n}{2}+1\right)}\times\sqrt{\prod_{i=1}^{n}\lambda_{i}},$
where $\lambda_{i}$ are the eigenvalues of the covariance matrix of spectrum representations of all synthetic models, $n$ is the number of dimensions, and $\Gamma$ denotes the gamma function. This formula determines the volume of an ellipsoid in n-dimensional Euclidean space.
Figure 12 displays the relative volume proportions obtained by removing each synthesis option compared to the total volume of all synthesized models. As seen in the table, removing any synthesis option reduces the volume of the spectrum representation, with proportions smaller than 1. This suggests that incorporating all variations enhances the diversity of synthesized models.
IV-C Ablation Study (EQ2)
TABLE IV: Ablation study on the number of seeds to train models based on each synthetic architecture.
@@ IV-C Ablation Study (EQ2)
TABLE V: Ablation study on the real dataset for fingerprinted image generation based-on synthetic models.
@@
TABLE VI: Ablation study on the input image for fingerprint extraction.
Input Image Verification Identification
Acc AUC Acc F1
RGB 79.22 87.42 60.33 59.71
Spectrum 83.09 90.76 73.94 74.44
Noise Spectrum 89.73 95.69 94.03 93.84
@@
In the ablation study section, we examine several key factors affecting the generalization performance of the fingerprint extractor. These factors include: the types of synthetic models and real images used for creating fingerprinted images, the selection of input signals extracted from these fingerprinted images for training the fingerprint extractor, and the implementation of a sliding window strategy during the inference phase. This comprehensive evaluation helps in understanding how each component contributes to the effectiveness of the fingerprint extraction. Note that in this section, the fingerprint extractor is trained only on synthetic models and evaluated on all the models in Table I, without including any real generative models in training. Our default configurations within the tables are highlighted in gray.
Synthetic Models. To evaluate the impact of individual synthesis options of synthetic models on the generalization of the trained fingerprint extractor in model verification and identification tasks, we conducted an ablation study. The results, presented in Table III, demonstrate that omitting any synthesis option reduces the generalization capabilities in model identification and verification tasks. Notably, the removal of the upsampling and activation layers significantly affects generalization performance. This observation aligns with the Frechet Frequency Distance scores listed in Table II, where the absence of activation and upsampling variation results in the highest scores, indicating the lowest fidelity of synthetic models.
@@
Real Dataset. In Table V, we evaluate how the choice of real datasets used to generate fingerprinted images with synthetic models affects the generalization performance of the trained fingerprint extractor. We consider three options for constructing the real image pool: the CelebA dataset, which exclusively contains face images; the LSUN dataset, featuring 20 different types of semantic images; and a composite of both datasets. The results indicate that the combination of the CelebA and LSUN datasets, which offers the richest semantic diversity, yields the best generalization performance. This phenomenon may be because the increasing the semantic diversity of fingerprinted images helps the fingerprint extractor focus on semantic-agnostic fingerprint features.
Input Signal. In Table VI, we assess the impact of different input signals extracted from fingerprinted images on the training of the fingerprint extractor. We consider three types of inputs: the RGB image directly, the DFT spectrum of the RGB image (Spectrum), and the DFT spectrum of the noise extracted from the RGB image (Noise Spectrum). The noise is extracted using a denoising process that primarily removes high-level semantic details from the image. As demonstrated in the table, the choice of input signal significantly affects the generalization ability of the fingerprint extractor. The ”Noise Spectrum” input yields the best performance, followed by ”Spectrum,” with the direct RGB image showing the least effectiveness. This outcome suggests that the Noise Spectrum most effectively enhances the high-frequency fingerprint information within the image, making the fingerprint features more extractable and distinct.
@@
Classical Models
Emerging Models
(a) POSE
(b) Ours (Only-Syn)
(c) POSE
(d) Ours (Only-Syn)
Figure 13: Compare existing method POSE [10] and our method in attributing classical and emerging models. POSE is trained on samples from classical models. As depicted in the figure, it effectively extracts distinct fingerprint patterns from these seen models (a), while struggles to generalize this capability to unseen, emerging models (c). In contrast, our method, despite not being trained on any of these testing models, demonstrates robust generalization abilities on the two groups of models (b,d).
IV-D Comparison with Existing Methods (EQ3)
Model Identification In Table VIII, we evaluate and compare our methods—Ours (Only-Syn) and Ours (Fine-tune)—against existing methods in the 1:N model identification scenario. The results show that most existing methods achieve high identification accuracy on classical models, which are seen in their training. However, their performance significantly declines when tested on unseen emerging models, dropping below 50%. In contrast, our method ”Ours (Only-Syn)” demonstrates strong generalization capabilities. Despite being trained only on synthesized models and not having exposure to real generative models, it achieves over 90% identification accuracy on the two sets of models. To align more closely with existing methods, we introduce ”Ours (Fine-tune),” which is further fine-tuned using classical model images. This adjustment improves the identification accuracy to 99% on classical models, with a modest enhancement on emerging models. These results underscore that our fingerprint extractor, by leveraging synthetic models, exhibits superior generalization ability across real-world generative models due to its exposure to a broader model fingerprint variations. Figure 13 visualizing the extracted fingerprint embeddings also indicates this.

### v3

@@ III-A4 Normalization Effects
III-A5 Effects of Different Generative Blocks
The discussion above focuses on the influence of basic components. Moving forward, it is important to consider how different generative blocks throughout the entire generation process contribute to the final spectral pattern on the generate image. Generative models typically consist of multiple generative blocks, with upsampling layers playing a crucial role as connectors between these blocks, doubling the resolution of the input feature maps. As mentioned above, commonly used upsampling layers, such as bilinear and nearest-neighbor interpolation, inherently possess a low-pass filtering effect. This property leads them to attenuate high-frequency components from prior layers. Given that existing research [6] suggesting that model fingerprints predominantly reside in the high-frequency components, it becomes apparent that the later blocks in a model’s architecture could more significantly influence the model fingerprints on the output image as the high-frequency patterns from earlier layers could be filtered out. To empirically test this hypothesis, we designed an experiment depicted in Figure 3, consisting of three key steps:
@@ III-A5 Effects of Different Generative Blocks
Figure 3(d,e) shows the confusion matrices under Testing I and Testing II of two experiments. We use two ProGAN models for Model 1 and Model 2. In (d), we constructed these models using nearest-neighbor upsampling layers. In (e), the models utilize bilinear upsampling layers. As shown in the Figure, under Testing I, the classifier achieves high accuracy, indicated by the high diagonal values in the confusion matrix. However, after swapping the last few blocks in Testing II, the confusion matrix reveals a notable reversal in model attribution results. Specifically, when swapping the last two blocks, the reversal accuracy reaches 0.9. These shifted results suggest that the most distinguishable spectral patterns for differentiating models are primarily generated by the last two blocks of the models. In contrast, the patterns left by earlier blocks appear to become less distinct as they progress through the generative process.
@@ III-B Model Synthesis Strategy
Overall, our model synthesis strategy is illustrated in Figure 4, the structure of the synthetic model could be viewed as a shallow auto-encoder, composed of $K$ downsampling blocks and $K$ upsampling blocks. Based on the above discussion, $K$ can selected as a small value, such as one or two. Each downsampling block use a fixed architecture, with a pooling layer to downscale the input resolution by half, and two convolutional layers to increase the feature dimension. The feature output from the downsampling blocks is then sent into the upsampling blocks, which share common components with the generative blocks in standard generative models but offer various architectural options.
In summary, the architecture of synthetic models is defined by the options $\{K,L,U,A,N,S\}$.
  • $K$ refers to the number of downsampling/upsampling blocks within the synthetic model, which can be one or two.
  • $L$ represents the number of convolution layers per block, which can be one or two.
@@
By varying these configurations, we can get $2\times 2\times 3\times 4\times 3\times 2=288$ different architectures within the construction space. For each architecture, we train $M$ models with different training seeds, creating $M$ distinct models with different parameters per architecture. $M$ is set to 20 in our experiments.
@@
We recognize that speed is a key factor in model synthesis. In our approach, we simplify the objective to focus on reconstruction, where the generative neural network is specifically tasked with minimizing the reconstruction loss of input images. This method is not only straightforward to implement but also allows for rapid training. We leave the exploration of other training approaches to future work. We set a constraint on the minimum reconstruction residual to $\eta=0.005$ to limit the amount of artifacts in the output images. Ultimately, the total number of synthetic models generated is 5760, allowing for a wide simulation of a wide range of potential generative models with diverse architectures and parameters.
III-C Effectiveness of Model Synthesis
We evaluate the effectiveness of our model synthesis strategy in terms of Fidelity and Efficiency. Fidelity assesses how well the synthetic models replicate the spectrum/fingerprint patterns of real-world generative models. Efficiency examines the speed of training on a large number of synthetic models.
Fidelity. Currently, there is no off-the-shelf tool to quantitatively measure the gap between synthetic and real models in fingerprint distributions. To address this gap, we introduce a new metric called Frechet Frequency Distance (FFD), inspired by the Frechet Inception Distance (FID) [27]. The choice of Frechet Distance is made due to its practical utility and widespread acceptance for comparing distributions, especially in contexts where these the mean and covariance effectively captures differences in the average content and the variation of the data. Different from FID that is used to assess the fidelity of generated images by comparing the distribution distance of inception features, FFD assess the fidelity of the fingerprints of synthetic models by comparing the distribution distance of spectrum features. The FFD is calculated through two primary steps:
Models All w/o K w/o L w/o S
FFD 3.30 3.57 3.41 3.51
Models w/o U w/o A w/o N w/o seed
FFD 9.81 4.46 3.02 3.39
  • (Step I) Spectrum Pattern Representation: We extract the spectrum pattern of generated images to represent the fingerprint pattern of each synthesized and real models. As depicted in Figure 5(a), for each model, we initially generate $N$ (e.g., 100) images [^4] We provide ablation study on the value of $N$ in Appendix -C., and then, following [13], we enhance the spectrum patterns by applying a noise extractor [28] to remove semantic contents from these images. The resultant noise images are then used to compute a reduced 1D power spectrum [17] by Azimuthal Integration (AI) over radial frequencies on the 2D Fourier spectrum. We calculate the averaged spectrum from these $N$ images to serve as the spectrum pattern representation for each model.
  • (Step II) Frechet Distance Calculation: As illustrated in Figure 5(b), we first compute the mean and covariance of the spectrum representations for all real-world models in Table III, which covers the mainstream generative models ranging from GAN, VAE, Flow and Diffusion. Then in the same way, we calculate the mean and covariance of the spectrum representations for synthetic models in Section 4. The Fréchet distance is then calculated between these distributions. Specifically, for the mean $\mu_{r}$ and covariance $\Sigma_{r}$ of real distribution, and the mean $\mu_{s}$ and covariance $\Sigma_{s}$ of synthetic distribution, their distance is given by:
    $\text{FFD}=\left\|\mu_{r}-\mu_{s}\right\|^{2}+\operatorname{Tr}\left(\Sigma_{r}+\Sigma_{s}-2\left(\Sigma_{r}\Sigma_{s}\right)^{1/2}\right),$ (10)
In Table II, we evaluate the Frechet Frequency Distance of synthetic models against real models, varying synthesis choices. The variations in the architecture include: “w/o K” and “w/o L,” indicating the construction with a single down/up block and a single layer, respectively; “w/o S,” using only one type of sequence — normalization and activation after convolution; “w/o U,” utilizing only bilinear interpolation for upsampling; and “w/o A” and “w/o N,” which denote the absence of activation and normalization layers, respectively. We also consider not using diverse seeds for each architecture, denoted as “w/o seed”. The results displayed in the table indicate that reducing most synthesis options tends to increase the Frechet Frequency Distance (FFD), pointing to a poorer alignment between the distributions. Notably, eliminating the diversity in activation functions and upsampling methods results in the most significant increases in FFD. This may be because these factors, when varied, mostly increase the diversity of synthetic models.
@@ III-C Effectiveness of Model Synthesis
Efficiency. We measured the time required to train these models. For models with single generative block ($K=1$), the averaged required training time per model is 53 seconds, while those with $K=2$ required 113 seconds each. In total, training all the synthetic models takes 133 hours on a single 3090 GPU. This is considerably more efficient than the time investment needed for training state-of-the-art generative models. For example, it takes 14 days 22 hours on single V100 GPUs to train a StyleGAN [14] model that generate with a 256$\times$256 resolution [^5] https://github.com/NVlabs/stylegan.
III-D Training the Fingerprint Extractor
@@
$I_{k}=M_{k}(I_{\text{real}})$ (11)
To reveal the underlying fingerprint patterns inherent in the images, we follow a common preprocessing technique [29, 30] to strip away low-frequency semantic content through a denoiser by [28]. The extracted noises are then transformed using a Discrete Fourier Transform to facilitate the extraction of their spectrum properties. The spectrum obtained from this transformation are fed into the fingerprint extractor, which is built upon the ResNet50 architecture.
$h_{k}$ $=f_{\operatorname{fp\_extractor}}\left({\mathcal{F}}(I_{k}^{\prime})\right)$ (12)
@@
To enhance the model’s performance, we combine classification and metric loss to enhance the discrimination of learned fingerprint embedding. The classification head employs a multi-class cross-entropy (CE) loss function, which enable the fingerprint extractor to accurately classify the originating synthetic model of each input image. The projection head is followed by a triplet loss introduced in [31], which encourages generated images from the same generator to be as similar as possible in the feature space while promoting a larger dissimilarity between images from different generators.
$\mathcal{L}=\mathcal{L}_{\text{CE}}(h_{k},y_{k})+\mathcal{L}_{\text{Triplet}}(h_{a},h_{p},h_{n})$ (14)
where $y_{k}$ is the true label for $k$-th synthetic model. $h_{a}$, $h_{p}$, and $h_{n}$ are the anchor, positive, and negative fingerprint embedding from a triplet set. Positive and negative mean the embedding is from the same model with the anchor embedding or not.
@@ IV Experiments
EQ4 How similar are synthetic models to real models in terms of their 2D spectrum visualizations?
@@ IV-A Experimental Details
Testing Models. Although our analysis and synthesis strategy primarily focuses on CNN-based generative models, our evaluation includes a broader scope of models as shown in Table III. In addition to CNN-based models, we also assess the generalization performance on Transformer-based models, which share many common components with CNN-based models. And for CNN-based category, we include a variety of models such as CNN, VAE, Flow, and Diffusion models. All models in the CNN-based and Transformer-based category are unconditional models, which are trained on the dataset in the “Image Source” column with different resolutions. Furthermore, the last row of the table encompasses state-of-the-art Text2Image models such as Stable Diffusion v1.5, Stable Diffusion v2.1, Stable Diffusion XL, Glide, DalleE-mini, DalleE-2, and DalleE-3. These models are not restricted to specific domains, and diverse in architecture type, training dataset, and resolution. Although these models are primarily for multimodal text-to-image generation, they share structural similarities with GAN-based unimodal generators. For example, the foundational architecture of the Stable Diffusion v1 series employs a U-Net structure, where its upsampling blocks are also built using standard components such as normalization, activation, convolution, and upsampling layers. Furthermore, as highlighted in  [13, 32, 33], no generator to date is entirely artifact-free, with frequency-domain artifacts still evident even in state-of-the-art multimodal generative models.
TABLE III: Models highlighted in gray are considered classical models and are used to train the existing methods being compared. The remaining models are categorized as emerging models and have not been seen by the existing methods.
@@
Testing Scenarios. Similar to the face identification and verification problems encountered in the field of face recognition, we consider utilize the fingerprint extractor in two model attribution problems, the 1:1 model verification and 1:N model identification problem, as shown in Figure 7. Settings and experimental setups for the two problems are illustrated below:
@@
  • Model Verification: the 1:1 model verification problem aims to verify whether two generated images are from the same model or not. Model verification is performed by comparing the cosine similarity of the extracted fingerprint. To assess the performance, we generate 1,000 images for each model under test. Then we randomly selected 10,000 pairs out of these images, which consist of 5,000 negative pairs (images from different models) and 5,000 positive pairs (images from the same model). The evaluation employs metrics commonly used in face verification tasks: accuracy and the Area Under the Receiver Operating Characteristic Curve (AUC). Accuracy measures the correct verification of pairs as originating from the same or different models. We report the best accuracy across all similarity thresholds to assess the overall effectiveness of the verification system. The AUC provides a comprehensive measure of performance across all possible threshold levels, reflecting the trade-off between the true positive rate and the false positive rate.
@@
  • Model Identification: the 1:N model identification problem aims to identify the specific generative model used to create a given image among N models. N is the total number of models in the model gallery. To assess the performance of model identification, we generate 1,000 images for each model under test. Out of these, 10 images are sampled to create the fingerprint for each model, while the rest serve as test samples. Each model’s fingerprint is collected to form the model gallery. We compare the test sample with each fingerprint in the gallery and classify it into the class with the highest cosine similarity, which corresponds to the closest fingerprint distance. Performance is evaluated using classification accuracy and the F1 score.
@@
Compared methods. To compare with existing methods, we derive two variants of our method: 1) Ours (Only-Syn), which is only trained on synthetic models without having seen any real-world models. 2) Ours (Fine-tune), which fine-tune the fingerprint extractor trained on synthetic models for one epoch using samples of models highlighted in gray in Table III.
@@
We compare against seven existing model attribution methods. They are Marra et al. [5], Yu et al. [6], DNA-Det [8], RepMIX [9], POSE [10], and Abady et al. [11]. We train them on models in gray in Table III. Although the works in [5, 6, 8, 9] are proposed to evaluate in a closed-set scenario, their the layers before the classification head can be adapted as a fingerprint extractor for use in a zero-shot model attribution context. This adaptation is implemented to evaluate the model identification and verification problems.
IV-B Ablation Study (EQ1)
In the ablation study section, we examine several key factors affecting the generalization performance of the fingerprint extractor. These factors include: the types of synthetic models and real images used for creating fingerprinted images, and the implementation of a sliding window strategy during the inference phase. This comprehensive evaluation helps in understanding how each component contributes to the effectiveness of the fingerprint extraction. Our default configurations within the tables are highlighted in gray. In this section, we evaluate on the Only-Syn version of our method, without including any real generative models in training.
TABLE IV: Ablation study on the architecture of synthetic models.
@@ IV-B Ablation Study (EQ1)
TABLE V: Ablation study on the number of seeds to train models based on each synthetic architecture.
@@
TABLE VI: Ablation study on the real dataset for fingerprinted image generation based-on synthetic models.
@@
Figure 8: (Left) Totaling training time for synthetic models with a maximum generative block count of 1, 2, and 3, corresponding to 2880, 5760, and 8640 models, respectively. (Right) The model verification accuracy by the fingerprint extractor trained on synthetic models with a maximum block count from 1 to 3.
Synthetic Options. To evaluate the impact of individual synthesis options of synthetic models on the generalization of the trained fingerprint extractor in model verification and identification tasks, we conducted an ablation study. The results, presented in Table IV, demonstrate that omitting any synthesis option reduces the generalization capabilities in model identification and verification tasks. Notably, the removal of the upsampling and activation layers significantly affects generalization performance. This observation aligns with the Frechet Frequency Distance scores listed in Table II, where the absence of activation and upsampling variation results in the highest scores, indicating the lowest fidelity of synthetic models.
Depth of Synthetic Models. We also discuss how depth of synthetic models influences the attribution performance. In Figure 8, we plot the total training time required for synthetic models with a maximum block count ranging from 1 to 3. For models with a single generative block, the averaged training time for all models is 42 GPU hours, involving models with 2 generative blocks, the training time is 133 hours, and increase to 393 hours when adding models with 3 generative blocks. Meanwhile, we observe an improvement in model verification accuracy when increasing the block count to 2, but no significant performance increase with 3 blocks. This indicates that using synthetic models with a maximum block count of 2 achieves the best trade-off between synthesis computation costs and model attribution performance.
@@
Real Dataset. In Table VI, we evaluate how the choice of real datasets used to generate fingerprinted images with synthetic models affects the generalization performance of the trained fingerprint extractor. We consider three options for constructing the real image pool: the CelebA dataset, which exclusively contains face images; the LSUN dataset, featuring 20 different types of semantic images; and a composite of both datasets. The results indicate that the combination of the CelebA and LSUN datasets, which offers the richest semantic diversity, yields the best generalization performance. This phenomenon may be because the increasing the semantic diversity of fingerprinted images helps the fingerprint extractor focus on semantic-agnostic fingerprint features.
@@
IV-C Comparison with Existing Methods (EQ2)
Model Identification In Table VIII, we evaluate and compare our methods—Ours (Only-Syn) and Ours (Fine-tune)—against existing methods in the 1:N model identification scenario. The results show that most existing methods achieve high identification accuracy on classical models, which are seen in their training. However, their performance significantly declines when tested on unseen emerging models, dropping below 50%. In contrast, our method “Ours (Only-Syn)” demonstrates strong generalization capabilities. Despite being trained only on synthesized models and not having exposure to real generative models, it achieves over 90% identification accuracy on the two sets of models. To align more closely with existing methods, we introduce “Ours (Fine-tune),” which is further fine-tuned using classical model images. This adjustment improves the identification accuracy to 99% on classical models, with a modest enhancement on emerging models. These results underscore that our fingerprint extractor, by leveraging synthetic models, exhibits superior generalization ability across real-world generative models due to its exposure to a broader model fingerprint variations.